Using the shrinkage allowances shown below, how much extra should be added to the mold dimension to get a cast dimension of 2.5 feet for a cast iron part?

Dimension

Shrinkage allowance (in/ft)

Up to 2 feet

0.125

2 to 4 feet

0.105

Over 4 feet

0.083
Answer #1

Material : Cast Iron

Cast dimensions : 2.5 feet

The acceptable shrinkage allowance for 2 feet to 4 feet dimensions is 0.105 in per feet

So the allowance for 1 inch is    

Then the final mold dimension for a cast of 2.5 feet dimension must be

= 2.5 + (0.00875*5)+(0.105*2)

= 2.5 +(0.04375+0.21)

=2.5 +0.25375

=2.75375 or 2.75 feet

thus the extra dimension that need to be added is 0.25375